OpenStudy (anonymous):

square root cscx-2=0

OpenStudy (anonymous):

Is the square root over everything?$$\sqrt{\csc x-2}=0\\\csc x-2=0\\...$$or did you mean:$$\sqrt{\csc x}-2=0\\\sqrt{\csc x}=2\\...$$